OpenStudy (anonymous):

7, 21,63, 189,... Use the formula for the sum of the first n terms of a geometric system to solve this exercise. Find the sum of the first 11 terms.

OpenStudy (anonymous):

looks like \(r=3\) so you have \[7+7\times 3+7\times 3^2+...+7\times 3^{10}\] \[=7(1+3+3^2+...+3^{10}\] \[=7\frac{3^{11}-1}{3-1}\]

OpenStudy (anonymous):

7, 21, 63, 189, 567, 1701, 5103, 15309, 45927, 137781 = 206668

OpenStudy (anonymous):

you need eleven terms, not ten

OpenStudy (anonymous):

oh my bad i read wrong.

OpenStudy (anonymous):

620011 should be the total.